[clas_members] [New Logentry] RC Daily Summary February 1, 2023

bueltman at jlab.org bueltman at jlab.org
Wed Feb 1 10:55:02 EST 2023

Logentry Text:
Day Shift:
The solenoid was successfully ramped to full field early and the 15 mm empty target cell used for target/beam position alignment runs.
>From 12:00 to 13:00 the target was changed to 20 mm carbon.
Carbon data taking with 4 nA beam (small adjustment of raster offset toward end of shift).

Evening Shift:
Carbon data taking with 4 nA until about 20:30 when the beam went off.
22:30 Solenoid isolation vacuum alarm, turbo pump stopped working. Denny came in to fix the problem.

Night Shift:
2:00 Turbo fore pump temporarily replaced and hall ready for beam.
Data taking on carbon until 6:30 when the beam stopped for Maintenance Day (beam was unavailable for about 1 hour during that time).

After completion of target change to ND3, polarize target positive and take production data with 4 nA.
Hall Lockup scheduled for 15:00.

We will have a RC meeting today at 2:15 PM.

RC Meeting Coordinates:
Meeting ID: 160 867 4253
Passcode: 786506


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4121710
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20230201/9bd9584d/attachment.html>

More information about the clas_members mailing list